OpenStudy (anonymous):

4 by the power of 2 -2(3x5+1) I have no clue what to do with the answer I got from the parenthesis since you are supposed to do that first.......

OpenStudy (templeguy):

the answer is -16

OpenStudy (irishboy123):

i wish i could read minds too :p

OpenStudy (anonymous):

Thank you soo much @templeguy

OpenStudy (anonymous):

@IrishBoy123 sorry if I was confusing I really don't explain things very well soooo

OpenStudy (freckles):

I'm sorry what is going on in that problem? are you saying: \[4^{2-2(3x^5+1)}\]

OpenStudy (freckles):

oh is that one x a multiplication sign

OpenStudy (freckles):

\[4^2-2(3 \times 5+1)\] after thinking about it I think you mean this

OpenStudy (freckles):

order of operations say to perform what is in the grouping symbols first

OpenStudy (freckles):

in the grouping symbols you have 3(5)+1

OpenStudy (freckles):

you have multiplication and addition there the multiplication comes first then the addition

OpenStudy (freckles):

so 3(5)+1=?

OpenStudy (anonymous):

\[4^{2}-2(3x5+1)\]

OpenStudy (freckles):

oh so that is an x and not a multiplication symbol

OpenStudy (freckles):

are you suppose to evaluate the expression for some x value?

OpenStudy (anonymous):

@freckles no it's just a multiplication symbol

OpenStudy (freckles):

so I had it right a sec ago \[4^2-2(3 \times 5+1)\]

OpenStudy (freckles):

order of operations: 1) grouping symbols 2) exponents 3) mult/div as occurs left to right 3) add/sub as occurs left to right

OpenStudy (freckles):

looking at 1) we do have grouping symbols contained in the grouping symbols is 3(5)+1 we are to do multiplication before addition so you have to compute 3(5) and then add 1 on to that to get the answer for 3(5)+1

OpenStudy (freckles):

and you know 3(5)+1 =?

OpenStudy (templeguy):

Answer these five questions for five metals

OpenStudy (anonymous):

16 @freckles

OpenStudy (freckles):

right so we have \[4^2-2(3 \times 5+1) \\ 4^2-2(15+1) \\ 4^2-2(16)\]

OpenStudy (freckles):

now we have exponents subtraction and multiplication

OpenStudy (freckles):

according to the order of operations we are to take care of our exponenets first

OpenStudy (freckles):

and remember 4^2 means 4 times 4

OpenStudy (irishboy123):

@Haley123450 :-))

OpenStudy (freckles):

\[4^2-2(3 \times 5+1) \\ 4^2-2(15+1) \\ 4^2-2(16) \\ 16-2(16)\] now only two operations remain subtraction and multiplication you are to do the multiplication first then the subtraction you should get @templeguy 's answer after completion

OpenStudy (anonymous):

I didn't know what to do with the other half because I had 16 but I didn't know if I as supposed to add it multiply it or what. So it was kinda just sitting there

OpenStudy (anonymous):

And I totally agree

OpenStudy (freckles):

16-2(16) yes this means we need to multiply 2 and 16 which gives us 32 and so you have 16-32 you can also write this as -32+16 or -(32-16) -(16) -16

OpenStudy (freckles):

anyways i hope you feel better about this problem